40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 6 14:36:15 2018, gautam, Update, LSC, TRX clipping 
    Reply  Thu Jun 7 16:15:03 2018, gautam, Update, LSC, TRX clipping 
Message ID: 13920     Entry time: Wed Jun 6 14:36:15 2018     Reply to this: 13927
Author: gautam 
Type: Update 
Category: LSC 
Subject: TRX clipping 

For some time now, I've been puzzled by the unreliability of the ASS_X dither alignment servo. Leaving the servo on, TRX often begins to decay to a lower value, and even after freezing the dither at the maximum TRX values, I can manually align the mirrors to increase TRX. We have suspected some kind of clipping in the TRX path that is responsible for this behaviour. Today I decided to investigate this a bit further. To have the arm locked and to inspect the beam, we have to change the locking trigger - TRX is what is normally used, but I misaligned the Y arm completely, and used AS110 as a trigger instead. There is some strangeness in the triggering topology, but this deserves a separate elog.

Once the arm was locked (and relocks using the AS110 trigger in the event of an unlock), I was able to trace the beampath on the EX table with an IR card. The TRX beam is rather large and weak, so it is hard to see, but as best as I can tell, the only real danger of clipping (or perhaps the beam is already clipped) is on the final steering mirror before the beam hits the (Thorlabs) PD. Steve/Pooja are working on getting a photo of this, and will upload it here shortly. Options to mitigate this:

  1. Use the harmonic separator to steer the beam lower, and center it on the 1" steering mirror. However, this could possibly lead to clipping on some of the upstream lenses.
  2. Raise the height of the 1" steering mirror by 0.25". However, this would require a custom 3/4" dia post height or some shims, which I am not sure is in line with our optomechanic mounting practises.
  3. Use a 2" mirror instead of a 1" mirror.

The EX QPD has stopped working since the Acromag install. If it were working, we wouldn't have to rely on the alternate triggering with AS110 and instead just use the QPD as TRX, while we debug the Thorlabs PD path.

ELOG V3.1.3-